### v1

# MUSCLE: A Model Update Strategy for Compatible LLM Evolution

Abstract
Large Language Models (LLMs) are frequently updated due to data or architecture changes to improve their performance. When updating models, developers often focus on increasing overall performance metrics with less emphasis on being compatible with previous model versions. However, users often build a mental model Bansal et al. (2019) of the functionality and capabilities of a particular ML model they are interacting with. They have to adapt their mental model with every update – a draining task that can lead to user dissatisfaction. In practice, fine-tuned downstream task adapters rely on pretrained LLM base models. When these base models are updated, these user-facing downstream task models experience instance regression or negative flips –previously correct instances are now predicted incorrectly. This happens even when the downstream task training procedures remain identical. Our work aims to provide seamless model updates to a user in two ways. First, we provide evaluation metrics for a notion of compatibility to prior model versions, specifically for generative tasks but also applicable for discriminative tasks. We observe regression and inconsistencies between different model versions on a diverse set of tasks and model updates. Second, we propose a training strategy to minimize the number of inconsistencies in model updates, involving training of a compatibility model that can enhance task fine-tuned language models. We reduce negative flips – instances where a prior model version was correct, but a new model incorrect – by up to 40% from Llama 1 to Llama 2.
MUSCLE: A Model Update Strategy for Compatible LLM Evolution
Jessica Echterhoffthanks:   Research conducted during internship at Apple. Corresponding authors: jechterh@ucsd.edu,    mpouransari@apple.com Fartash Faghri Raviteja Vemulapalli UC San Diego Apple Apple Ting-Yao Hu Chun-Liang Li Oncel Tuzel Hadi Pouransari Apple Apple Apple Apple
1 Introduction
Large Language Models (LLMs) are often pretrained on large-scale corpora to obtain a base model that has general world knowledge. These base models are typically evaluated using a suite of benchmarks that mostly focus on zero/few-shot performance and in-context learning capabilities. Training these models is expensive and only a few organizations have access to the resources needed for this. Hence, to enable various user-facing applications such as summarization, chatbots, code assistants, and question answering, practitioners often adapt pretrained base models by training task-specific parameter-efficient adapters using downstream task datasets.
Figure 1: A real example of a model update that introduces negative flips, where previously correct predictions become incorrect. With our model update strategy using a compatibility adapter approach, we can enhance compatibility with the previous model while maintaining the performance gain (measured by the ROUGE-1 score for the summarization task in this figure) of the model update.
Several scenarios drive updates to the base model. For example, improved training strategy, advances in LLM architectures or increasing model context length Touvron et al. (2023a), the availability of additional or higher quality datasets Gunasekar et al. (2023), the expansion of model vocabulary (e.g., to support multilingual or multimodal models), or simply training for a longer period  Biderman et al. (2023). When a base model is updated, the associated task adapters need to be retrained to have meaningful downstream models. Hence, in the rest of the paper, we use the term model update to refer to an update to a downstream task model, which includes updating the base model and retraining the task adapter.
When a model is updated, each data sample can be categorized into one of the four groups depicted by four quadrants shown in Fig. 2. The new model could result in a negative flip (quadrant 4) for many samples (Yan et al., 2021) even when it has a better overall performance when compared to the previous model. A real example of a negative flip caused by a model update is shown in Fig. 1 for diagloug summarizatoin task. This regression behavior could confuse the user driving down the user satisfaction (Sakai, 2022). Furthermore, the importance of regression testing has been recently studied for evolving LLMs used through APIs (Ma et al., 2023).
Many existing works focus on increasing positive flips (quadrant 2) while reducing negative flips (quadrant 4) Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). However, yhey neglect prediction inconsistencies in the scenarios where both model versions are incorrect (quadrant 3). For example, Träuble et al. (2021) assumes the cost of flips from one incorrect class to another incorrect class to be zero. We argue that there is value in being consistent when both models are incorrect. A user may have developed coping strategies on how to interact with a model when it is incorrect; therefore, inconsistencies in mistakes can lead to user dissatisfaction.
Previous works have mainly addressed the regression challenge for classification tasks Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). In this work, we systematically study the problem of downstream model compatibility (reducing negative flips and being consistent when making mistakes) using various downstream tasks (both discriminative and generative) and different base models.
Following is a summary of our contributions:
  • We formulate the compatibility problem when updating LLMs. We consider a setup where base LLMs are updated independently and study model compatibility and performance through parameter-efficient task-specific fine-tuning.
  • We extend notion of model compatibility from discriminative to generative tasks, and propose regression/compatibility metrics that consider similarities in model behavior after an update, going beyond the negative flip rate metric.
  • We investigate model compatibility for different update scenarios using open-weights models and find significant regression across various tasks.
  • We propose learning a compatibility adapter to align model versions and minimize regression. We demonstrate up to 40% reduction in negative flip rate (e.g. for Llama 1 to Llama 2 update in language understanding) and reduction in model inconsistency for other downstream tasks such as summarization, math reasoning, and commonsense question-answering.
Figure 2: Four possibilities arise for each sample when a model is updated. Quadrants 2 and 4 show positive and negative flips, respectively. A model update with more instances in quadrant 2 and fewer in quadrant 4 indicates more performance gain and less regression. Furthermore, quadrant 3 corresponds to instances where both models are incorrect. Encouraging similarity between the old and new models in this case (i.e., making the same mistakes) results in a more seamless model update from the user’s perspective.
2 Related Work
2.1 Measuring Regression Errors
Classification
Yan et al. (2021) introduce the negative flip rate (NFR) to evaluate model compatibility for classification tasks. NFR calculates the fraction of instances that were previously correct but are now incorrect with the new model. Other metrics include backward trust compatibility (BTC) (Srivastava et al., 2020), the ratio of instances that the new model predicts correctly among all instances the old model predicts correctly. Ultimately, these metrics have been shown to measure similar statistics Sakai (2022). Matsuno and Sakuma (2023) propose backward compatibility with a conditional distribution, which computes the ratio at which the accuracy of the conditional distribution of the new model is equal to or higher than that of the old model. Cai et al. (2022) introduce the negative flip impact for graph NLP tasks, taking into account the negative flips and the overall error rate of the model. All of these metrics are limited to the evaliation of discriminative classification tasks.
2.2 Reducing Update Regression
Model Ensembles
Prior work found that model ensembles reduce model update regression. This can be attributed to the reduction of variance by ensembling, as every single model may capture the training data from a distinct aspect (Yan et al., 2021; Xie et al., 2021). Other approaches include choosing the most centric model from the ensemble Xie et al. (2021), aligning two models’ uncertainties Li et al. (2023b), or using gating mechanisms Lai et al. (2023). To ensure model compatibility, previous work has also used model parts from the old model to infuse into the new one (Ran et al., 2023), with the limitation of both models being required at inference time. For limited use cases, Qin et al. (2023) have shown to re-use previously learned adapters when purely a data update was performed. These methods either introduce larger memory footprint by re-using old model parts, or are limited to (same-domain) data-updates and same models only.
Knowledge Distillation
Originally proposed for model compression Buciluǎ et al. (2006), a (smaller) student model is trained to mimic a (larger) teacher model. By treating the old model as the teacher and the new model as the student, knowledge distillation has been shown to reduce model update regression in vision and language discriminative tasks Yan et al. (2021); Xie et al. (2021); Schumann et al. (2023); Jaeckle et al. (2023); Zhang et al. (2021); Shen et al. (2020); Ramanujan et al. (2022). Shen et al. (2020) propose a distillation-based influence loss to align new model representations with those of the old model. Similarly, Ramanujan et al. (2022); Jaeckle et al. (2023) apply distillation after a learned transformation module. Schumann et al. (2023) propose weight interpolation between the old and the new model, Zhao et al. (2022) suggest matching old and new model distributions, Yan et al. (2021) distill from an ensemble, and Caciolai et al. (2023) recommend using focal distillation. However, none of these approaches evaluate their approaches on generative tasks or require both the old and new models to be in memory.
3 Problem Formulation
Setup
Training LLMs from scratch for every downstream task is expensive. Hence, we follow the common setup of finetuning a pre-trained base LLM to multiple downstream tasks with task-specific LoRA adapters (Hu et al., 2021).
Let $\mathcal{M}^{\text{base}}_{i}$ denote the $i^{th}$ version of the base LLM with parameters $\theta_{i}$. We adapt $\mathcal{M}^{\text{base}}_{i}$ to a downstream task $\mathcal{T}$ using an adapter $\mathcal{A}_{i}^{\mathcal{T}}$ to obtain a downstream model $\mathcal{M}_{i}^{\mathcal{T}}$ with weights $\theta^{\mathcal{T}}_{i}=\theta_{i}+\Delta_{i}^{\mathcal{T}}$, where $\Delta_{i}^{\mathcal{T}}$ denotes the weights of the task-specific adapter $\mathcal{A}_{i}^{\mathcal{T}}$ learned using the training data corresponding to task $\mathcal{T}$. When the base model is updated from $\mathcal{M}^{\text{base}}_{v1}$ to $\mathcal{M}^{\text{base}}_{v2}$, the task-specific adapters are re-trained for each downstream task. Hereafter, for simplicity of notation, we use $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ to refer to the task-adapted models $\mathcal{M}_{v1}^{\mathcal{T}}$ and $\mathcal{M}_{v2}^{\mathcal{T}}$, respectively, and explicitly mention the task $\mathcal{T}$ when needed.
Figure 3: When updating a model, regression on individual tokens and instances can arise. We use a masked approach to select tokens to be aligned with knowledge distillation either with the old version to remain consistent, or with the new task model to increase performance.
3.1 Backward Compatibility Metrics
A backward compatibility metric outputs a compatibility score based on two models, $\mathcal{M}_{v1}$ to $\mathcal{M}_{v2}$. Yan et al. (2021) propose negative flip rate (NFR) to measure regression in classification models over a dataset $\{x_{i},y_{i}\}_{i=1}^{N}$, where $y_{i}$ is the ground truth class for input $x_{i}$ for a particular task $\mathcal{T}$:
$\text{NF}(x_{i})\triangleq[\mathcal{M}_{v1}(x_{i})=y_{i}]\land[\mathcal{M}_{v2}(x_{i})\neq y_{i}]$
$\text{NFR}\triangleq\frac{1}{N}\sum_{i}^{N}\mathbf{1}[\text{NF}(x_{i})]$
Here $\mathbf{1}$ denotes the indicator function. This notion of regression is partly applicable for autoregressively trained tasks. LLM benchmarks Gao et al. (2023) calculate the likelihood of every possible choice in a multiple-choice question and choose the response with the highest likelihood. This evaluation can be used to indicate regression, similar to classification for tasks with multiple choices Zellers et al. (2019); Wang et al. (2019); Welbl et al. (2017).
Unobserved Inconsistencies
Inconsistencies do not only occur within positive or negative flips. For example, when the old model predicts class A, the new model class B, and the ground truth is class C. Similarly, if there are multiple ground truth options, a flip could occur within the ground truth options, as in quadrant 1 in Fig. 2. We argue that if we cannot produce a better prediction, we should at least stay consistent to the users expectations and propose metrics extended negative flips metrics for this use case.
Continuous Metrics
In generative tasks for language models, we do not necessarily have multiple choices for which we can predict if an instance regressed or not. Hence a metric incorporating multiple choices and then calculating negative flips is not applicable for continuous evaluation metrics. For example, for summarization tasks, which continuous metrics such as ROUGE (Lin, 2004) or BERT Score (Zhang et al., 2019) scores are often used for evaluation. Hence, we require a new metric to evaluate compatibility for generative tasks like summarization.
3.2 Extended Evaluation Metrics
To extend the possibility of evaluating model updates, we propose a suite of metrics that evaluate model updates specifically for generative tasks that may or may not have multiple options available (e.g summarization tasks), as well as extend measurement of inconsistent flips.
Accounting for Flips within Gain or Regression
To capture inconsistencies for instances where both old and new models are incorrect, we adapt the negative flip rate as follows when multiple choice options are available:
$\text{NF}_{mc}(x_{i})\triangleq$ $[\mathcal{M}_{v2}(x_{i})\neq y_{i}]\land$
$[\mathcal{M}_{v1}(x_{i})\neq\mathcal{M}_{v2}(x_{i})]$
$\text{NFR}_{mc}\triangleq\frac{1}{N}\sum_{i}^{N}\mathbf{1}[\text{NF}_{mc}(x_{i})]$
which includes the possibility that neither of the models give the correct answer, but a change in behavior occurs. Similarly, for multi-label tasks this notion can account for ground-truths that may have flipped during a model update when multiple grount truth options exist.
Smooth Gain and Regression
To add fine-grained and continuous metrics for generative tasks like summarization, we evaluate the expected performance gain and regression compared to the previous version, to enable a user-centric model update in accordance to the four quadrants of Fig. 4. Our general framework aims to be independent of the actual similarity metric, such that it can be chosen in accordance with any respective task of interest.
Given a similarity metric $S$, and model outputs $\mathcal{M}_{v1}(x_{i})$ and $\mathcal{M}_{v2}(x_{i})$ for an input $x_{i}$, the difference for a model update for a particular test instance $i$ is
$D(x_{i})\triangleq S(\mathcal{M}_{v2}(x_{i}),y_{i})-S(\mathcal{M}_{v1}(x_{i}),y_{i})$
acting as an indication of the distance of the two model outputs with respect to the grand truth. This gain per instance enables us to classify which model update compatibility quadrant each instance falls into. In practice, similarity metrics could be BERT Score Zhang et al. (2019), ROUGE Score Lin (2004), BLEU Score Papineni et al. (2002), or model-as-a-judge metrics Huang et al. (2024) depending on the use case and task.
To get an indication similar to positive and negative flip rate in generative tasks, we observe the distribution of instances with positive gain or negative regression:
$\widetilde{\text{PFR}}\triangleq\frac{1}{N}\sum_{i}^{N}\mathbf{1}[D(x_{i})>0]$
$\widetilde{\text{NFR}}\triangleq\frac{1}{N}\sum_{i}^{N}\mathbf{1}[D(x_{i})<0]$
To observe the magnitude of regression and gain, we can observe the expectation:
$m_{g}\triangleq\frac{1}{N\cdot\widetilde{\text{PFR}}}\sum_{i}^{N}D(x_{i})~{}\mathbf{1}[D(x_{i})>0]$
$m_{r}\triangleq\frac{1}{N\cdot\widetilde{\text{NFR}}}\sum_{i}^{N}|D(x_{i})|~{}\mathbf{1}[D(x_{i})<0]$
This gives an indication of the average magnitude of change in similarity metric when gain or regression occurs.
$\mathcal{M}^{\text{base}}_{v1}$ $\mathcal{M}^{\text{base}}_{v2}$ Update
Phi 1 Phi 1.5 Synthetic data, data selection Gunasekar et al. (2023); Li et al. (2023a)
Phi 1.5 Phi 2 Synthetic data, data selection, $\uparrow$ parameters Javaheripi et al. (2023)
Llama 1 Llama 2 More data, $\uparrow$ context size Touvron et al. (2023a, b)
Vicuna 1.3 Vicuna 1.5 Llama 1 $\rightarrow$ Llama 2, Instruction fine-tuning Zheng et al. (2024)
Table 1: We select a suite of models with varying update scenarios and parameter sizes and evaluate them on different LLM benchmark tasks. We select Llama and Vicuna models with 7B parameters.
Dataset Task
HellaSwag (Zellers et al., 2019)
Language Understanding
(Multiple-Choice)
PIQA (Bisk et al., 2019)
Commonsense Reasoning
(Binary-Choice)
GSM8k (Cobbe et al., 2021)
Grade School Math
(Exact Match)
SAMsum (Gliwa et al., 2019) Dialogue Summarization
Table 2: We tackle compatible model updates for different downstream tasks and datasets, including multiple-choice and generative tasks.
4 Knowledge Transfer
[…]
$\mathcal{L}_{KL}=\frac{1}{n}\sum_{i=1}^{n}KL(\sigma(z_{t,i}/T)\|\sigma(z_{s,i}/T))$
Most knowledge distillation works consider the distillation from a trained teacher to an untrained student Tian et al. (2022); Rajasegaran et al. (2020). Recent work (Roth et al., 2024) tackles the goal of knowledge transfer between pretrained student and teacher models while retaining student knowledge gained a priori, and show that standard knowledge distillation between pretrained models struggles to transfer knowledge without performance drops for many teacher-student pairs. Complementary to this work that focuses on performance and maintaining prior knowledge, we tackle compatibility to previous models through knowledge transfer.
4.1 Model Update Strategy for Compatible LLM Evolution (MUSCLE)
When the base model is updated, we aim to train a task-specific fine-tuned model, $\mathcal{M}^{C}_{v2}$, that has the accuracy benefits of $\mathcal{M}_{v2}$, but with most compatibility with $\mathcal{M}_{v1}$. We obtain $\mathcal{M}^{C}_{v2}$ by training a compatibility adapter applied to the base model $\mathcal{M}^{\text{base}}_{v2}$. We use knowledge from task-specific fine-tuned models $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ when training $\mathcal{M}^{C}_{v2}$. $\mathcal{M}_{v2}$ typically has increased prediction capabilities over $\mathcal{M}_{v1}$ (due to improvements in the base model), but $\mathcal{M}_{v1}$ has information on already correctly predicted tokens or instances that we want to minimize regression towards.
We initialize the compatibility adapter with the task-specific adapter of $\mathcal{M}_{v2}$, and further fine-tune it (using the task training dataset) by aligning the next token prediction to either $\mathcal{M}_{v1}$ or $\mathcal{M}_{v2}$. We define a masking (for individual tokens of a training sequence) following a simple heuristic depending on whether $\mathcal{M}^{C}_{v2}$ (the adapter being trained) predicts the correct tokens or not. If it does, we align to $\mathcal{M}_{v2}$ logits, otherwise, we align to $\mathcal{M}_{v1}$. The fine-tuning process of $\mathcal{M}^{C}_{v2}$ is depicted in Fig. 3. The fine-tuning loss to train the compatibility adapter, $\mathcal{L}_{comp}^{m}$, is defined below:
$m_{i}=\mathbf{1}[\text{argmax}~{}\sigma(z_{\mathcal{M}^{C}_{v2},i})\neq y_{i}]$ (1)
$\begin{split}a_{\mathcal{M}_{v1}}=&KL(\sigma(z_{\mathcal{M}_{v1},i}/T)\|\sigma(z_{\mathcal{M}^{C}_{v2},i}/T))\\ a_{\mathcal{M}_{v2}}=&KL(\sigma(z_{\mathcal{M}_{v2},i}/T)\|\sigma(z_{\mathcal{M}^{C}_{v2},i}/T))\\ \mathcal{L}_{comp}^{m}=&\frac{1}{n}\sum_{i=1}^{n}m_{i}\cdot a_{\mathcal{M}_{v1}}+(1-m_{i})\cdot a_{\mathcal{M}_{v2}}\end{split}$ (2)
When evaluating, we denote NFR as negative flip rate between $\mathcal{M}_{v1}$ $\mathcal{M}_{v2}$, $\text{NFR}_{c}$ as the observed negative flip rate between $\mathcal{M}_{v1}$ and our compatibility model $\mathcal{M}^{C}_{v2}$, and $\Delta\text{NFR}_{c}=\text{NFR}_{c}-\text{NFR}$.
$\mathcal{M}^{\text{base}}_{v1}$ $\mathcal{M}^{\text{base}}_{v2}$ $\text{acc}_{\mathcal{M}_{v1}}$ $\text{acc}_{\mathcal{M}_{v2}}$ $\text{acc}_{c}$ $\Delta\text{acc}_{c}\uparrow$ NFR $\Delta\text{NFR}_{c}\downarrow$ $\Delta_{\%}\text{NFR}_{c}\downarrow$
HellaSwag Llama 1 Llama 2 72.74 72.91 79.53 6.62 10.27 -4.17 -40.60
Phi 1.5 Phi 2 69.76 78.02 77.86 -0.16 3.09 0.24 7.77
Phi 1 Phi 1.5 32.38 69.76 69.86 0.10 2.47 -0.13 -5.26
Vicuna 1.3 Vicuna 1.5 72.19 71.54 78.10 6.56 10.48 -4.06 -38.74
PIQA Llama 1 Llama 2 74.86 74.76 79.27 4.51 11.59 -3.97 -34.25
Phi 1.5 Phi 2 74.65 78.78 78.94 0.16 6.20 -0.11 -1.77
Phi 1 Phi 1.5 59.03 74.65 74.70 0.05 7.78 -0.05 -0.64
Vicuna 1.3 Vicuna 1.5 74.70 75.52 78.89 3.37 9.90 -2.88 -29.09
Table 3: Compatible task adapter trained with MUSCLE (corresponding to metrics with suffix $c$) reduces negative flip rate. We see enhanced improvements for model updates that have smaller performance difference.
$\mathcal{M}^{\text{base}}_{v1}$ $\mathcal{M}^{\text{base}}_{v2}$ $\text{EM}_{\mathcal{M}_{v1}}$ $\text{EM}_{\mathcal{M}_{v2}}$ $\text{EM}_{c}$ $\Delta\text{EM}_{c}\uparrow$ NFR $\Delta\text{NFR}_{c}\downarrow$ $\Delta_{\%}\text{NFR}_{c}\downarrow$
Llama 1 Llama 2 24.45 33.09 36.66 3.57 8.49 -0.91 -10.72
Phi 1.5 Phi 2 30.02 48.18 50.68 2.50 5.88 -1.71 -29.08
Phi 1 Phi 1.5 3.41 30.02 26.99 -3.03 2.01 -0.04 -1.99
Vicuna 1.3 Vicuna 1.5 26.72 29.91 31.84 1.93 11.60 -0.99 -8.53
Table 4: GSM8K math evaluation with exact match (EM). For compatibility adapter (corresponding to metrics with suffix $c$), we observe decreased negative flip rate while mostly maintaining performance gains.
5 Experimental Setup
5.1 Model Update Assumptions
To analyse the impact of model updates we consider parameter efficient fine-tuned models using Low-Rank Adapters (LoRA) Hu et al. (2021). Compared to previous work on continuous learning and model updates Qin et al. (2022, 2023), we do not limit model updates to be produced by only data updates, but consider different kinds of updates shown in Table 1. We include updates due to data, increased parameters, or different training strategies.
More importantly, we include wide range of downstream tasks to evaluate model compatibility, including generative tasks as summarized in Table 1. For all tasks we learn a LoRA adapter autoregressively (next-token prediction). Some of these tasks can be evaluated with classification-like approaches, e.g., by choosing the maximum log-likelihood of the possible answers (Gao et al., 2023), such as commonsense reasoning tasks (Zellers et al., 2019; Bisk et al., 2019). For other tasks such as dialogue summarization (Gliwa et al., 2019), we cannot evaluate in a classification-like manner.
5.2 Task Adapter Training
For each task and each old model $\mathcal{M}_{v1}$ and new model $\mathcal{M}_{v2}$, we train a LoRA adapter on all linear layers with $r=128$ and $\alpha=256$. We use a 0.8/0.2 training/validation split for 10 epochs. We choose the model by cross-entropy validation loss.
5.3 Compatibility Adapter Training
We keep all hyper-parameters the same to task adapter training, but train compatibility adapter with the $\mathcal{L}_{comp}^{m}$ loss defined in Eq. 2. We analyze statistical significance of the results on a subset of compatibility adapter training for Phi 1 to Phi 1.5 updates on PIQA using 3 random seeds. We observe a standard deviation of accuracy of 0.0012.
We compare $\mathcal{L}_{comp}^{m}$ with different masking strategies $m$ in the ablation studies in Section 6.5, but find that the version introduced in Section 4.1 works best for most model updates. Further, for model updates with large performance gaps between $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ we find that an auxiliary cross-entropy loss enhances stability of training for PIQA and SAMSum from Phi 1 to Phi 1.5, and all Phi updates in GSM8k and Hellaswag.
Figure 4: When updating LLM models (e.g. Llama 1 $\rightarrow$ Llama 2), we observe negative flips in different tasks. The smaller the performance gap from an old model to a new model, the more negative flips we observe. When evaluating continuous metrics with absolute ROUGE-1 value for summarization on SAMsum Gliwa et al. (2019), we observe a large fraction of negative flips. We show the exact models analysed in Table 8.
6 Results
6.1 Negative Flips Occur in Model Updates
Fig. 4 shows that significant negative flips (up to more than 60%) exist for a variety of base model update scenarios and downstream tasks. We observe negative flips in updates within one model family (e.g. Llama/Vicuna, and Phi models). We find more negative flips for model updates with a smaller delta in performance gain. For generative tasks like SAMsum dialogue summarization, we observe a large number of negative flips, $\text{NFR}_{s}$, as continuous metrics are more sensitive to small changes when updated.
6.2 Reduced Negative Flips in Classification
Binary and multiple choice generative tasks can be evaluated with classification metrics, such as calculating the log likelihood of each answer option (we use Gao et al. (2023) for this use case). This can be interpreted as a “classical” sense of evaluating negative flips. In Table 3, we observe that MUSCLE decreases negative flips as shown when compared to regular model update without compatibility specific training ($\mathcal{M}_{v2}$). Specifically, we show reduction of NFR by 40% for the update of LLama 1 to Llama 2 and 39% for Vicuna 1.3 to Vicuna 1.5. For updates with large accuracy gap (for example Phi 1 to Phi 1.5), we observe a less strong enhancement with a negative flip rate reduction by 1-5%. In addition to reduction of negative flips we also observe an increased accuracy of up to 7% for Llama and Vicuna updates.
Similar to multiple-choice tasks, we can evaluate math questions in a classification-like approach to determine negative flips. In exact match (EM) evaluation we match the final result of the math question from the prediction to ground truth. Results are shown in Table 4. In this case, we observe that we can reduce the number of negative flips by 29% for Phi 1.5 to Phi 2. When the version 1 model is significantly less accurate (e.g., 3.4% for Phi 1), we observe a reduction in accuracy while only being able to decrease negative flips by 2%. For all other updates, MUSCLE increases expected-match accuracy while reducing negative flips.
Figure 5: Comparison of NFR vs NFRmc metrics to evaluate inconsistency when updating LLMs for HellaSwag task.
6.3 Increased Consistent Behavior
When we cannot achieve a positive flip (switching from an incorrect to a correct answer), we might prefer to at least maintain consistent behavior to old model to avoid unexpected behavior for the user. We evaluate negative flips and inconsistency flips. For model updates that have a large performance gap and a small number of negative flips to begin with, we see limited reduction in inconsistency flips. However, we observe that we can reduce the inconsistency flips for model updates with smaller accuracy gap such as the updates for Llama and Vicuna (Fig. 5) on the HellaSwag dataset.
[…]
Llama 1 Llama 2 32.06 32.28 34.79 2.51 48.96 -7.81 -15.95 0.64 -1.13
Phi 1.5 Phi 2 37.53 36.15 40.69 4.54 54.70 -15.02 -27.46 0.24 -3.39
Phi 1 Phi 1.5 30.92 37.53 38.76 1.23 32.60 -5.74 -17.61 -0.38 -0.59
Vicuna 1.3 Vicuna 1.5 30.32 30.88 34.08 3.20 49.69 -10.98 -22.10 -0.02 -1.85
Table 5: For summarization (SAMsum) generative task we reduce regression of ROUGE-1 score (R1) up to 27%.
6.4 Reduced Regression in Generative Tasks
When evaluating generative tasks, such as summarization, we observe that we can reduce regression of ROUGE-1 score performance (Table 5). We can reduce negative flips by 18% for updates with weaker version 1 models (Phi 1 $\rightarrow$ Phi 1.5) and 22% for smaller model updates (Vicuna 1.3 $\rightarrow$ Vicuna 1.5) and 27% for Phi 1.5 $\rightarrow$ Phi 2. When looking at average gain of ROUGE 1 Score we see that we decrease ROUGE-1 regression by 1-3%, while maintaining gain.
$\Delta\text{NFR}_{c}$ $\do\Delta_{\%}\text{NFR}_{c}\do$ $\Delta\text{PFR}_{c}$ $\Delta\text{acc}_{c}$
$\mathcal{L}_{comp}^{\mathcal{M}^{C}_{v2}\neq y}$3.97 -34.25 0.54 4.51
$\mathcal{L}_{comp}^{\mathcal{M}_{v1}=y}$2.29 -19.76 -0.60 1.68
$a_{\mathcal{M}_{v1}}$2.99 -25.80 -0.82 2.17
CE+$\mathcal{L}_{comp}^{\mathcal{M}^{C}_{v2}\neq y}$3.10 -26.75 0.38 3.48
CE+$\mathcal{L}_{comp}^{LL_{L}}$2.12 -18.29 -0.49 1.63
CE+$\mathcal{L}_{comp}^{LL_{S}}$1.96 -16.91 0.05 2.01
Table 6: A model update from Llama 1 $\rightarrow$ Llama 2 on PIQA. Different ablations and masking strategies and their impact on negative flips, positive flips and accuracy improvement.
6.5 The Effect of Different Masking Strategies
In this section, we analyze different training and masking strategies to evaluate our design choices. On the PIQA dataset and model update Llama 1 $\rightarrow$ Llama 2, we compare MUSCLE with different masking strategies. Intuitively, likelihood based masking strategies can be useful for tasks that are evaluated with log-likelihoods (e.g. PIQA, HellaSwag) where multiple choices are available. We compare likelihood based masking for example on individual tokens, $m_{i}=LL_{L}=\mathbf{1}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]$ to check if the likelihood of the correct next token in the current model is smaller than the old model. Only in this case, we align to $\mathcal{M}_{v1}$, and align to $\mathcal{M}_{v2}$ for every other token. Alternatively, we can also compare the likelihood of the entire sequence $m_{i}=LL_{S}=\mathbf{1}[\sum_{i}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]]$, such that we mask all tokens per sequence to get instance-wise masking. In both cases, we see a reduction in negative flips. However, likelihood based masking requires auxiliary cross-entropy loss for stability.
Furthermore, aligning to the old model only if it is already correct with a mask ($m_{i}=\mathbf{1}[\sigma(z_{\mathcal{M}_{v2},i}))=y_{i}]$) or purely (no masking) with KL ($a_{\mathcal{M}_{v1}}$) also lead to a small reduction in negative flips. Aligning to the old model if the current prediction is incorrect, i.e., Eq. 1, leads to the best reduction in negative flips while providing the biggest accuracy gain and increased positive flips (Table 6). For this best-performing strategy, we see that including cross entropy loss does not lead to performance gains. This strategy has the additional advantage that it takes into account extended inconsistent flips like explained in Fig. 5.
7 Conclusion
In this work, we study the task-specific compatibility problem when updating LLMs. We show that LLM updates with different scenarios, e.g., changes in model architecture, optimization, or training dataset, exhibit significant negative flips– instances previously classified or generated correctly, and incorrectly after the model update. We extend negative flip metric for discriminative and, for the first time, generative tasks, and report results for various models and tasks.
We propose a novel method (MUSCLE) to train task-specific compatibility adapter when updating an old LLM to a new LLM to reduce negative flips while maintaining performance gain. Our proposed method does not require a modification to the base model training, and is only based on adapter training. Further, as opposed to previous works, the proposed solution does not require both versions of the model in memory to enable compatibility, which is often infeasible due to large size of LLMs.
We observe a mitigation of negative flips of 40% for multiple-choice type evaluations, and 27% for continuous summarization evaluation. We also show insights into model properties that facilitate transfer, finding that our alignment masking strategy provides best results with an additional benefit of mitigating inconsistent update behavior.
8 Limitations
In the current study we do not consider a model update that includes changes in tokenization and/or vocabulary size (e.g.Llama 2 Touvron et al. (2023b) to LLama 3 AI@Meta (2024)). Future work can explore compatible vocabulary mapping strategies before learning from prior model versions. Our results indicate that the utility of aligning to prior model versions is limited by the performance of the prior model version. For example, see the update from Phi 1 to Phi 1.5 in Table 4, where Phi 1 only has an accuracy of 3%. In this case, it is arguable if an alignment to $\mathcal{M}_{v1}$ is desired and if the strive for compatibility outweighs a possible performance drop.
9 Ethical Considerations and Risks
One potential risk of the proposed approach for compatible task-specific LLM updates is the transfer of potential biases from the old model, $\mathcal{M}_{v1}$, to the new model trained through knowledge transfer. We did not explore this aspect in the current study.
10 Acknowledgements
We would like to thank Rick Chang, Cheng-Yu Hsieh and Yen-Ju Lu for their help with the paper.
Appendix A Appendix
A.1 Training Hyperparameters and Evaluation
Hyperparameter Value
Epochs 10
Learning Rate 1e-4
Grad. acc. steps 8
LoRA $\alpha$ 156
LoRA rank 128
Dropout 0.0
Adapter Layers All linear
Warmup Steps 500
Weight decay 0.0
Table 7: Hyperparameters for the training setup of the task and compatibility adapters.
We show an overview of the design choices of our LoRA adapter for task and compatibility training in Table 7. We learn on the entire context and answer tokens during fine-tuning. We find that successfully training the compatibility adapter generally requires larger LoRA rank. Task adapters can be trained with lower rank, but also experience performance boost with higher rank. We use the training splits of the respective datasets where we use 80% of the training set as training data, and 20% as validation data for selecting the best model based on validation loss. We evaluate for selection after each epoch. All compatibility adapter models are initialized with the previously trained task adapter models for version 2. KL divergence temperature is set to 2. We use deepspeed for optimization Rasley et al. (2020).
For benchmark evaluation, we use LM evaluation harness Gao et al. (2023), where most of the benchmarks (GSM8k, HellaSwag, PIQA) are already defined, and we use them as-is. We add evaluation for SAMsum summarization evaluating ROUGE-1 score with a no repeat N-gram size of 2, and generation stop words of “Dialogue:”, “Summary:” (which are the key words for the context and answer behavior), “</s>” and double new lines. All of our tasks are based on the English language. All Experiments were run on NVIDIA A100 and H100 GPUs with a compute budget of 720x8 GPUh.
A.2 Model Update Evaluation
In the evaluation of regression for different model updates (Fig. 4), we consider the model pairs shown in Table 8.
$\mathcal{M}^{\text{base}}_{v1}$ $\mathcal{M}^{\text{base}}_{v2}$
Phi 1 Phi 1.5
Phi 1 Phi 2
Phi 1.5 Phi 2
Vicuna 1.3 Vicuna 1.5
Llama 1 Llama 2
Llama 1 Vicuna 1.5
Llama 1 Vicuna 1.3
Vicuna 1.3 Llama 2
Llama 2 Vicuna 1.5
Llama 2 Llama 3
Llama 1 Llama 3
Vicuna 1.3 Llama 3
Vicuna 1.5 Llama 3
Table 8: Model pairs used to analyse regression in model updates.
A.3 Downstream Tasks
Dataset Train Test
HellaSwag (Zellers et al., 2019) 40k 10k
PIQA (Bisk et al., 2019) 16k 3k
GSM8k (Cobbe et al., 2021) 7.5k 1.3k
SAMsum (Gliwa et al., 2019) 14.7k 819
Table 9: Dataset size for our experiments.
We describe the datasets used in our experiments. Each dataset poses unique challenges and targets different aspects of language understanding and reasoning.
GSM8K
The Grade School Math 8K (GSM8K) dataset designed to evaluate mathematical reasoning capabilities. It consists of grade school-level math word problems. These problems require the application of mathematical concepts and the ability to reason about quantities and operations in a textual format. The dataset is structured to test the performance of models on arithmetic, algebra, geometry, and statistics problems, reflecting a wide range of mathematical knowledge in education. We use this dataset as a representative case to test problem solving cases. We evaluate regression in this task with exact match, which is a strict evaluations on a fraction of produced tokens, but does not account for the reasoning process.
SAMSum
The SAMSum dataset consists of dialogue summaries designed to facilitate the evaluation of automatic conversational summarization models. It contains dialogue instances, paired with human-written summaries. These conversations mimic real-life scenarios to enable learning generate coherent and concise summaries. We use this dataset as a representative case to test language generation behavior. We evaluate regression in this task with ROUGE-1, which is a lenient evaluations on a fraction of produced tokens that are contained in the ground truth data.
HellaSwag
HellaSWAG is a dataset for assessing common sense reasoning and predictive text modeling. It builds on the SWAGZellers et al. (2018) dataset by providing more challenging distractors. HellaSWAG consists of multiple-choice scenarios where a model must predict the most plausible continuation among four options, focusing on everyday activities, scenarios, and interactions. We use this dataset as a representative case to test abilities that require not just linguistic understanding but also real-world knowledge and common sense reasoning. We evaluate regression in this task with log-likelihoods for the correct answer, as multiple choices are given.
PIQA
The Physical Interaction QA (PIQA) dataset tests the understanding of physical and causal interactions in the real world through textual descriptions. It contains scenarios that require reasoning about physical properties, actions, and outcomes. Each scenario is presented as a question with two possible solutions, where the model must choose the most physically plausible one. We use this dataset as a representative case for evaluating models on tasks that require an understanding of the physical world and its governing principles. We evaluate regression in this task with log-likelihoods for the correct answer, as different choices are given.